# V2XPnP: Vehicle-to-Everything Spatio-Temporal Fusion for Multi-Agent Perception and Prediction

### v2

@@ C Cooperative Temporal Perception Task
In addition to the end-to-end perception and prediction task, the sequential nature of our V2XPnP-Sequential dataset facilitates other temporal tasks, including temporal perception and traditional prediction tasks. In this section, we introduce the cooperative temporal perception task and present benchmark results on the V2XPnP-Sequential dataset. Details on the traditional prediction task are provided in Sec. D.
@@ C.2 Benchmark Methods
For benchmarking, we adapt our end-to-end model, V2XPnP, by removing the prediction head, resulting in a model only for temporal perception. Various V2X fusion strategies are evaluated in this framework, as detailed in Tab. S2. Moreover, we provide another baseline FaF, which adopts a combination of 2D and 3D convolutions for temporal fusion. FaF further integrates with the F-Cooper intermediate fusion method and early fusion method for V2X fusion comparison. We also provide the results of the No Temp model, which excludes temporal fusion and is evaluated using both F-Cooper and early fusion methods. Model parameters and experimental setups for this task are consistent with those used for the end-to-end cooperative perception and prediction task.
@@ D.1 Problem Formulation
V2XPnP sequential dataset also supports the traditional prediction task. Compared to end-to-end models, which directly infer future states of objects from perception data, the traditional prediction task forecasts their future trajectories from historical trajectories. The cooperative prediction task is formulated as: given the map and the historical trajectories of all detected objects obtained from the ego agent and other agents (e.g., CAVs and infrastructure units) within the communication range of the ego agent, the objective is to predict future trajectories of these detected objects.
@@ D.2 Benchmark Methods
To investigate the influence of perception results on prediction tasks, we provide two types of input for the prediction models: 1) Ground-truth historical trajectories of surrounding objects; 2) Perception-based historical trajectories generated by the upstream perception module. The first one is the common setting for the traditional trajectory prediction task, assuming full availability of accurate historical trajectories for prediction. However, it ignores real-world challenges such as occlusions and cumulative errors introduced by separate modules. To address this limitation and enable a more realistic evaluation, we designed the second setting, where CAVs can only derive the historical trajectories from the perception results and thus the perception uncertainty can propagate to the downstream prediction. Notably, regardless of the input type, the prediction model is trained using the complete future trajectory dataset aggregated from all agents.
@@ D.3 Benchmark Results
The experimental results, summarized in Tab. S3, compare traditional prediction under three input settings: ground truth trajectories, perception without fusion, and perception with late fusion. The results indicate that as perception improves—from no fusion to late fusion—the prediction performance correspondingly increases. When the environment is fully observable, the task simplifies to the traditional prediction setup, achieving the best overall performance for both detection and prediction. A significant drop in performance is observed for perception-based prediction, highlighting the critical dependency of predictive tasks on perception accuracy. Moreover, the Attention predictor shows better robustness compared to the LSTM baseline under noisy perception inputs, thanks to the attention module for complex interaction feature capturing. We anticipate that this temporal prediction task will inspire further exploration of perception-based prediction approaches.
Figure S1: Examples of interaction scenarios from the V2XPnP-Sequential dataset. The dataset multi-agent perception perspectives and captures diverse interaction behaviors among ten object classes in dense urban traffic environments.
@@ E.1 Dataset Visualization
Our V2XPnP-Sequential dataset provides two sensor sequences (LiDAR and camera) collected in dense urban environments, capturing diverse interactive behaviors over time. Fig. S1 illustrates two representative interaction scenarios in our dataset, presenting LiDAR and camera data from multiple agents at two key timestamps. The main intersection objects pair have been annotated with red and yellow blocks in different agents’ views.
@@ E.2 Data Acquisition
Coordinate System. Our V2XPnP-Sequential dataset encompasses three coordinate systems: the LiDAR coordinate system, the camera coordinate system, and the map coordinate system. Each agent - vehicle or infrastructure - maintains its own local LiDAR and camera coordinate systems. The global map coordinate system serves as the reference for all annotations and maps. The transformation from each agent’s local LiDAR coordinate to the map coordinate in each frame is achieved with the GPS/IMU data and the offline PCD map. We also conduct the 3D-2D calibration for LiDAR and camera, as shown in Fig. 4(b).
@@ E.3 Data Annotation and Processing
Trajectory Generation. In addition to perception data, the dataset provides a ground-truth trajectory dataset derived from the fused perception data of all agents, capturing the trajectories of objects across all frames. This trajectory dataset is primarily utilized in traditional prediction tasks, which assume all history trajectories are observable to the ego agent. However, this assumption ignores the fact that the trajectories obtained from onboard sensors are incomplete due to occlusion and limited perception range, and no specific datasets are designed to support this task. To support research in prediction with real-world sensor constraints, we provide a trajectory retrieve module in the V2XPnP-Sequential dataset to return observable trajectories of surrounding objects based on their actual visibility relationships.
Map Generation. The HD map generation involves two stages: point cloud (PCD) map generation and vector map generation. (1) To generate the PCD map, each LiDAR frame from the CAVs is preprocessed to remove dynamic objects, retaining only static elements essential for mapping. Then, a Normal Distributions Transform (NDT) scan-matching algorithm is employed to compute the relative transformation between consecutive frames, forming the basis of the LiDAR odometry. We also incorporate translation and heading information obtained from the vehicle’s GPS/IMU system, integrating them through a Kalman filter to refine the pose estimation, mitigating the drift from the error accumulation in LiDAR data. Finally, the LiDAR sequences are fused to form the PCD map across all collection areas. (2) The aggregated PCD map is imported into RoadRunner [7] to generate vector maps. Road geometry is inferred and annotated based on intensity variations visualized by distinct color mappings within RoadRunner, and all the semantic attribution is annotated based on the collected camera data, such as road type (e.g., driving, sidewalk, and parking) and line type (e.g., solid and broken yellow line combination and solid white line). Finally, the generated maps are exported in the OpenDRIVE (Xodr) format and converted to Waymo map format [9], ensuring compatibility with downstream applications.
@@ E.4 Dataset Analysis
Tab. S4 presents the comparison of the V2XPnP-Sequential dataset with existing driving datasets. Our dataset tracks an average of 136 objects per scene, recording high-density and complex traffic scenarios. Furthermore, the dataset’s extensive map and trajectory data further enhance its utility in cooperative perception and prediction research across all collaboration modes. The data distribution of ten object classes is shown in Fig. 4(d). The dataset covers 24 intersections of varying types, including roundabouts, T-junctions, and crossroads, as shown in Fig. S2. Notably, many collection areas have a significant gradient, which can facilitate the detection and prediction research in diverse terrain conditions.
@@ E.5 Dataset Privacy Protection
The V2XPnP-Sequential dataset is designed with stringent privacy safeguards to ensure the anonymity of individuals and vehicles. Trajectory data only include object IDs and positions, eliminating the possibility of tracking specific entities. All perception data undergoes privacy-preserving processing, with LiDAR annotations retaining only essential attributes such as object ID, agent type, and bounding box pose. Additionally, all image data has been anonymized, with human faces and other potentially sensitive details obscured or removed.

### v3

@@ C Cooperative Temporal Perception Task
In addition to the end-to-end perception and prediction task, the sequential nature of our V2XPnP-Sequential dataset facilitates other temporal tasks, including temporal perception and traditional prediction tasks. In this section, we introduce the cooperative temporal perception task and present benchmark results on the V2XPnP-Sequential dataset. Details on the traditional prediction task are provided in Sec.˜D.
@@ C.2 Benchmark Methods
For benchmarking, we adapt our end-to-end model, V2XPnP, by removing the prediction head, resulting in a model only for temporal perception. Various V2X fusion strategies are evaluated in this framework, as detailed in Tab.˜S2. Moreover, we provide another baseline FaF, which adopts a combination of 2D and 3D convolutions for temporal fusion. FaF further integrates with the F-Cooper intermediate fusion method and early fusion method for V2X fusion comparison. We also provide the results of the No Temp model, which excludes temporal fusion and is evaluated using both F-Cooper and early fusion methods. Model parameters and experimental setups for this task are consistent with those used for the end-to-end cooperative perception and prediction task.
@@ D.1 Problem Formulation
V2XPnP sequential dataset also supports the traditional prediction task. Compared to end-to-end models, which directly infer future states of objects from perception data, the traditional prediction task forecasts their future trajectories from historical trajectories. The cooperative prediction task is formulated as follows: given the map and the historical trajectories of all detected objects obtained from the ego agent and other agents (e.g., CAVs and infrastructure units) within the communication range of the ego agent, the objective is to predict future trajectories of these detected objects.
@@ D.2 Benchmark Methods
To investigate the influence of perception results on prediction tasks, we provide two types of input for the prediction models: 1) Ground-truth historical trajectories of surrounding objects; 2) Perception-based historical trajectories generated by the upstream perception module. The first one is the common setting for the traditional trajectory prediction task, assuming full availability of accurate historical trajectories for prediction. However, it ignores real-world challenges such as occlusions and cumulative errors introduced by separate modules. To address this limitation and enable a more realistic evaluation, we designed the second setting, where CAVs can only derive the historical trajectories from the perception results, and thus the perception uncertainty can propagate to the downstream prediction. Notably, regardless of the input type, the prediction model is trained using the complete future trajectory dataset aggregated from all agents.
@@ D.3 Benchmark Results
The experimental results, summarized in Tab.˜S3, compare traditional prediction under three input settings: ground truth trajectories, perception without fusion, and perception with late fusion. The results indicate that as perception improves—from no fusion to late fusion—the prediction performance correspondingly increases. When the environment is fully observable, the task simplifies to the traditional prediction setup, achieving the best overall performance for both detection and prediction. A significant drop in performance is observed for perception-based prediction, highlighting the critical dependency of predictive tasks on perception accuracy. Moreover, the Attention predictor shows better robustness compared to the LSTM baseline under noisy perception inputs, thanks to the attention module for complex interaction feature capturing. We anticipate that this temporal prediction task will inspire further exploration of perception-based prediction approaches.
Figure S1: Examples of interaction scenarios from the V2XPnP-Sequential dataset. The dataset provides multi-agent perception perspectives and captures diverse interaction behaviors among ten object classes in dense urban traffic environments.
@@ E.1 Dataset Visualization
Our V2XPnP-Sequential dataset provides two sensor sequences (LiDAR and camera) collected in dense urban environments, capturing diverse interactive behaviors over time. Fig.˜S1 illustrates two representative interaction scenarios in our dataset, presenting LiDAR and camera data from multiple agents at two key timestamps. The main intersection objects pair has been annotated with red and yellow blocks in different agents’ views.
@@ E.2 Data Acquisition
Coordinate System. Our V2XPnP-Sequential dataset encompasses three coordinate systems: the LiDAR coordinate system, the camera coordinate system, and the map coordinate system. Each agent (vehicle or infrastructure) maintains its own local LiDAR and camera coordinate systems. The global map coordinate system serves as the reference for all annotations and maps. The transformation from each agent’s local LiDAR coordinate to the map coordinate in each frame is achieved with the GPS/IMU data and the offline PCD map. We also conduct the 3D-2D calibration for LiDAR and camera, as shown in Fig. 4(b).
@@ E.3 Data Annotation and Processing
Trajectory Generation. In addition to perception data, the dataset provides a ground-truth trajectory dataset derived from the fused perception data of all agents, capturing the trajectories of objects across all frames. This trajectory dataset is primarily utilized in traditional prediction tasks, which assume all history trajectories are observable to the ego agent. However, this assumption ignores the fact that the trajectories obtained from onboard sensors are incomplete due to occlusion and limited perception range, and no specific datasets are designed to support this task. To support research in prediction with real-world sensor constraints, we provide a trajectory retrieval module in the V2XPnP-Sequential dataset to return observable trajectories of surrounding objects based on their actual visibility relationships.
Map Generation. The HD map generation involves two stages: point cloud (PCD) map generation and vector map generation. (1) To generate the PCD map, each LiDAR frame from the CAVs is preprocessed to remove dynamic objects, retaining only static elements essential for mapping. Then, a Normal Distribution Transform (NDT) scan-matching algorithm is employed to compute the relative transformation between consecutive frames, forming the basis of the LiDAR odometry. We also incorporate translation and heading information obtained from the vehicle’s GPS/IMU system, integrating them through a Kalman filter to refine the pose estimation, mitigating the drift from the error accumulation in LiDAR data. Finally, the LiDAR sequences are fused to form the PCD map across all collection areas. (2) The aggregated PCD map is imported into RoadRunner [7] to generate vector maps. Road geometry is inferred and annotated based on intensity variations visualized by distinct color mappings within RoadRunner, and all the semantic attribution is annotated based on the collected camera data, such as road type (e.g., driving, sidewalk, and parking) and line type (e.g., solid and broken yellow line combination and solid white line). Finally, the generated maps are exported in the OpenDRIVE (Xodr) format and converted to Waymo map format [9], ensuring compatibility with downstream applications.
@@ E.4 Dataset Analysis
Tab.˜S4 presents the comparison of the V2XPnP-Sequential dataset with existing driving datasets. Our dataset tracks an average of 136 objects per scene, recording high-density and complex traffic scenarios. Furthermore, the dataset’s extensive map and trajectory data further enhance its utility in cooperative perception and prediction research across all collaboration modes. The data distribution of ten object classes is shown in Fig. 4(d). The dataset covers 24 intersections of varying types, including roundabouts, T-junctions, and crossroads, as shown in Fig.˜S2. Notably, many collection areas have a significant gradient, which can facilitate the detection and prediction research in diverse terrain conditions.
@@ E.5 Dataset Privacy Protection
The V2XPnP-Sequential dataset is designed with stringent privacy safeguards to ensure the anonymity of individuals and vehicles. Trajectory data only includes object IDs and positions, eliminating the possibility of tracking specific entities. All perception data undergoes privacy-preserving processing, with LiDAR annotations retaining only essential attributes such as object ID, agent type, and bounding box pose. Additionally, all image data has been anonymized, with human faces and other potentially sensitive details obscured or removed.